OpenStudy (anonymous):

find the equation of each line in slope intercept form: A verticle line through the point (2,4)

OpenStudy (rajee_sam):

do you know the graphs of vertical and horizontal lines. How they will look like?

OpenStudy (anonymous):

yes i do

OpenStudy (anonymous):



OpenStudy (anonymous):

quandrant 1

OpenStudy (anonymous):

with that drawn, i am trying to tell you the equation of the line is x=2...since it is passing through x-axis at point 2.

OpenStudy (anonymous):

waht about his one throught the point (6,1) with slope=1/3 i got y=1/3x-1 is tht correct

OpenStudy (jdoe0001):

stick those (6,1) values in the point-slope form, along with the 1/3, and expand

OpenStudy (jdoe0001):

that is, \( y-(6) = \cfrac{1}{3}(x-1) \)

OpenStudy (anonymous):

since the slope is 1/3 plug it in y=mx+c \[y=\frac{ 1 }{ 3 }x+c\] plug in your x and y-coordinates to find the value of c...

OpenStudy (anonymous):

what do you get for c??

OpenStudy (anonymous):

i dont know am looking at htes lesson again sorry am slow

OpenStudy (anonymous):

what is the x-coordinate in (6,1)?? and what is the y-coordinate??

OpenStudy (anonymous):

do u have a bright storm account?

OpenStudy (anonymous):

no....